OpenStudy (anonymous):

How do I factor 1-x^2

OpenStudy (anonymous):

difference of 2 square factoring (1-x)(1+x)

OpenStudy (anonymous):

even isf the x^2 is negative ?

OpenStudy (anonymous):

yes, because 1 and x^2 are both perfect squares and they are separated by a - sign. Multiply the factored form that I gave you and you'll get what you started with.

OpenStudy (anonymous):

Difference of squares (a^2-b^2) = (a+b)(a-b)

OpenStudy (anonymous):

\[(a+b)(a-b) = a(a-b) + b(a-b) = a^2 -ab +ba -b^2 = a^2-b^2\]

OpenStudy (anonymous):

x^2-1 is (x+1)(x-1)

OpenStudy (anonymous):

and 1^2 - x^2 is?

OpenStudy (anonymous):

U see? 1 = 1^2

OpenStudy (anonymous):

dhatraditya How does that make any sence can you just tell me what the answer is ? please.

OpenStudy (anonymous):

U have been told the answer 4 times already.

OpenStudy (anonymous):

ab and ba cancel each other. that is why (a+b)(a-b) = a^2-b^2 sorry, I cannot give you the answer, you have to work toward it yourself. I can help guide you to it.